OpenStudy (anonymous):

Please explain. 4x-2(1-x)=-38 4x-2+2x=-38 6x-2=-38 6x=-36 x=-6 On the second line when you move the +2x over dont you add +4 buy -2? But then I get 2 not 6! Please help.

OpenStudy (anonymous):

They just added the 4x to the 2x, both are on the left side

OpenStudy (jackellyn):

4x-2(1-x)=-38 4x -2(1) -(2)(-x) =-38 4x -2 +2x =-38 6x -2 =-38 6x =-36 x=-6

OpenStudy (anonymous):

OHHHHH ok I understand! I feel so stupid!

OpenStudy (anonymous):

Hehe, we've all been there!